​Multiple-choice questions each have four possible answers left parenthesis a comma b comma c comma d right parenthesis​, one of which is correct. Assume that you guess the answers to three such questions. a. Use the multiplication rule to find ​P(WCC​), where C denotes a correct answer and W denotes a wrong answer. ​P(WCC​)equals nothing ​(Type an exact​ answer.) b. Beginning with WCC​, make a complete list of the different possible arrangements of two correct answers and one wrong answer​, then find the probability for each entry in the list. ​P(WCC​)minussee above ​P(CCW​)equals nothing ​P(CWC​)equals nothing ​(Type exact​ answers.) c. Based on the preceding​ results, what is the probability of getting exactly two correct answers when three guesses are​ made? nothing ​(Type an exact​ answer.)
Answer #1

Answer:

P(C) = 1/4 and P(W) = 3/4

a) P(WCC) = (1/4)*(3/4)*(1/4) = 3/64 = 0.046875

b) P(WCC) = 3/64 = 0.046875

    P(CWC) = 3/64 = 0.046875

c) P(Two correct answers when three guesses are made) = 3*3/64 = 9/64 = 0.140625